OpenStudy (anonymous):

x^2y^2-x^4y / xy divisor 2 or more term long division

OpenStudy (anonymous):

\[(1 + x^6)/(x^4 - x^2 +1)\]

OpenStudy (anonymous):

is that how you do it or the answer?

OpenStudy (kinggeorge):

\[{x^2y^2-x^4y \over xy}={xy \over xy} \cdot (xy-x^3)=xy-x^3\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (kinggeorge):

If you wanted to find \[{x^2y^2-x^4y \over xy}\]then yes, that is the answer.

OpenStudy (anonymous):

thanks